# When Federated Learning Meets Quantum Computing: Survey and Research Opportunities

### v2

@@ 7.3 Quantum Secure Multi-Party Computation
In conventional FL systems, challenges such as gradient inversion attacks, which enable the reconstruction of private gradients, are prevalent. The authors in [107] addressed this issue by utilizing quantum states to encode gradients, thereby securing multi-party contributions to the global model from adversaries. The proposed framework incorporates protocols that use private inner-product estimation and incremental learning, collectively allowing contributions from each party without exposing individual gradients. These quantum-based techniques secure model aggregation among multiple parties by exploiting quantum phenomena, such as superposition and entanglement. In another work [109], a quantum secure aggregation scheme is proposed to facilitate secure multi-party contributions by aggregating local models using quantum secret sharing, where participants encode model gradients into qubits and transmit them via Greenberger–Horne–Zeilinger (GHZ) states [120]. Quantum entanglement detects eavesdropping, ensuring data confidentiality. Further, in [121], a hybrid-quantum-key based secure FL system is designed for multi-party applications, particularly in the medical field. This system uses quantum key distribution to securely share parameters and hybrid quantum keys to protect data exchanges and model contributions from external attacks. Combining quantum cryptographic techniques with classical methods enables secure multi-party communication.
@@
Analyzing $\mathrm{QUE}$ for secure multi‐party computation schemes indicates near‐maximal qubit utilization across protocols. The inner‐product encoding in [107] uses a four‐qubit, three‐layer circuit ($\mathrm{Q}_{alloc}=3$, $L=3$) obtaining $\mathrm{QUE}=1.0$ and thus adopts synchronous aggregation to leverage full‐batch quantum metric estimation; the GHZ‐based secret sharing in [109] runs on a three‐qubit, two‐layer design ($\mathrm{Q}_{alloc}=3$, $L=2$) with $\mathrm{QUE}=1.0$ and similarly favors synchronous updates to maintain entanglement integrity; and the quantum‐key DFL in [121] employs a five‐qubit, four‐layer ansatz ($\mathrm{Q}_{alloc}=5$, $L=4$) to achieve a high $\mathrm{QUE}\approx 1.0$ paired with hybrid classical–quantum fusion to balance key‐distribution overhead with model convergence. Despite their efficient parameter‐gate ratios and aligned model strategy choices, these SMPC frameworks contend with significant limitations. Firstly, they assume ideal entanglement distribution and noise‐free communication, omit analysis under non‐IID data splits, and lack adaptive mechanisms for client dropout or key‐renewal scheduling. Future work should explore decoherence‐aware aggregation protocols, resilient key management strategies in dynamic networks, and benchmark suites that jointly measure privacy, QUE, and end‐to‐end latency in dynamic QFL environments.
@@ 8 Miscellaneous
This section aims to elaborate on the practical usage and implementation of QFL, while discussing studies that apply it on real-life datasets that demonstrate QFL’s true potential. It also discusses the trainability and learnability aspects of QFL to further discuss challenges and opportunities .
@@ 8.1 Real-life Applications of QFL
Numerous other studies have also explored QFL to enhance model performance and address the computational demands in diverse application domains. Fig. 10 shows the broad application domains where QFL research has shown its footprints.
In healthcare contexts, federated quantum designs often emphasize data confidentiality. One dementia-classification method [122] employs QKD for encrypting parameter updates and safeguarding sensitive MRI data across hospital nodes. Despite a 1% increment in computational overhead, classification accuracy matched the baseline CNN (roughly 94%) while guaranteeing that eavesdroppers cannot easily invert gradients. Another proposal targets secure intrusion detection in consumer networks [123], achieving up to 3–4% gains in detection rates on complex traffic features when harnessing small-scale QNNs and homomorphic gradient aggregations. Addressing data heterogeneity is a central challenge in large-scale QFL. The Dynamic Aggregation QFL (DAQFL) approach [124] introduces an adaptive weighting scheme based on local accuracy improvements. By assigning higher weights to nodes with better QNN performance, the global model converges more reliably under non-IID splits. In healthcare IoT benchmarks, DAQFL achieves a 2–6% accuracy gain over baseline quantum federated averaging when dealing with skewed or long-tailed data distributions. QFL has also shown efficacy for sensitive clinical tasks such as pain-level detection. In [125], ECG signals are processed with a quantum-transfer-learning pipeline, integrated within a federated scheme to preserve patient privacy. Compared to conventional CNN-based methods, the quantum model improved classification accuracy by nearly 2%, reaching approximately 94.8% for multi-tier pain labeling. It also minimized data-exchange overhead by localizing heavier computations on quantum backends.
@@
Moving into automotive applications, QFSM [126] addresses speech-emotion recognition in vehicular networks by employing quantum minimal-gated recurrent units. This design shortens the gating operations and cuts training convergence time by roughly 10–15%. Field tests in a 5G IoV context report up to 2% higher accuracy than classical gated recurrent units on large-scale SER corpora, while also offering improved noise robustness in fluctuating channel conditions. Vehicular networks have also been extended into metaverse-based coordination. A quantum-enhanced FL framework [127] uses hierarchical clustering of vehicular nodes, where QNN states are compressed via quantum principal-component analysis. Adaptive reinforcement scheduling then modulates local updates, ensuring resource balancing across immersive metaverse tasks. Simulation results reflect around 15% lower communication overhead compared to standard FL, retaining near-baseline accuracy for trajectory prediction in high-mobility environments. A related thread investigates quantum FL in edge or drone-assisted scenarios. One surveillance solution [128] allows UAVs to transmit partial circuit parameters, dynamically adjusting QNN layer depths based on instantaneous channel quality. This reduces the communication load by nearly 20% while retaining top-1 accuracy comparable to a static-depth model on an aerial license-plate dataset. Such dynamic QNN pruning illustrates an essential synergy between quantum resource limits and federated training demands in bandwidth-constrained networks.
Another notable approach, ESQFL [129], integrates digital twin frameworks for real-time power-grid stability analysis. By linking quantum FL updates to digital replicas of physical bus nodes, the global model can detect incipient voltage collapse and dynamically schedule corrective actions. Shapley value calculations explain each node’s contribution in an interpretable manner, while differential privacy ensures data confidentiality. Grid simulations showed a 4% improvement in detecting instability events and a 10% boost in overall reliability.
Beyond the model-training process, some works explore hybrid optimization for FL resource management. One Benders’ decomposition algorithm [130] splits scheduling of workers and continuous learning-rate decisions into subproblems, solved alternately on a quantum annealer and a classical solver. At moderate scales, the method reduces runtime by up to 70% versus purely classical MILP solutions, indicating promise for large-scale FL scheduling. A further refinement involves quantum gate customization. Entanglement-controlled QFL [131] employs specialized universal gates to fine-tune the entanglement entropy in multi-depth QNNs. This strategy mitigates the barren-plateau problem by balancing inter-qubit interference. Fidelity-based distillation regularizers are added to stabilize training. Experiments indicate consistent convergence on non-iid data with minimal gradient vanishing.
@@ 8.2 Trainability and Learnability of QFL
Federated quantum models combine the expressive power of parameterized quantum circuits with the decentralized optimization of federated learning. Yet they inherit two intertwined theoretical challenges: ensuring effective optimization (trainability) and guaranteeing robust generalization (learnability).
As circuit depth $L$ or qubit count $Q$ increases, gradient magnitudes can vanish exponentially, a phenomenon known as the barren plateau [66, 39]. In QFL, each client’s local update may fall below machine noise, forcing the server to perform many more communication rounds to recover signal [52]. Empirical results on hardware-efficient ansatz show that limiting depth to $L\leq 4$ sustains gradients above numeric precision for up to 16 qubits, reducing required rounds by 30–50% compared to deeper designs [132]. Quantifying convergence, recent studies measure the number of gradient‐descent steps or federated rounds needed to reach a target accuracy. For example, FQNGD achieves 90% of its asymptotic accuracy within 20 rounds on standard benchmarks, compared to 40 rounds for FedAvg with the same ansatz [52]. This halving of communication cost stems from quantum natural gradient descent, which preconditions updates by the inverse quantum Fisher information matrix, enhancing per‐round progress up to 2× on ideal simulators [52]. Additional acceleration techniques further improve trainability. Layerwise training optimizes one ansatz block at a time and preserves gradient signal by avoiding global barren plateaus. Alternatively, smart initialization reuses parameters from smaller or related circuits to jump‐start optimization, cutting VQA iterations by up to 60% in QAOA settings [90, 133]. Together, these methods yield hybrid protocols that converge in fewer than 25 rounds even under moderate noise, making QFL practical on NISQ hardware. Further research could explore tensor networks for QFL and efficient algorithms with optimized ansatzs, beyond basic QAOA approaches.
On the other hand, learnability examines how well a QFL model trained on local data samples extends to unseen inputs. Classical FL theory derives generalization bounds via VC‐dimension and algorithmic stability; quantum analogues are emerging. Shallow, low‐entanglement circuits behave like kernel machines with controlled Rademacher complexity, suggesting that $\mathcal{O}(\log N)$ qubits suffice to generalize on $N$ training examples [40]. Yet over‐parameterized quantum models exhibit a double descent curve i.e., after an interpolation threshold, additional parameters can improve generalization under appropriate regularization [134, 135].
@@
where $n_{\rm classical}$, $n_{\rm quantum}$ are local sample sizes and $C_{\rm classical},C_{\rm quantum}$ depend on circuit expressivity [136]. Stability‐based bounds further relate the sensitivity of the global model to perturbations in client updates, with quantum algorithmic stability results implying that small‐norm updates preserve generalization under noisy mixing [137].
Fully characterizing the federated barren‐plateau frontier requires mapping how depth $L$, width $Q$, and client heterogeneity jointly determine gradient norms and convergence rates. On the learnability side, deriving optimal VC‐dimension and Rademacher‐complexity bounds for hybrid quantum–classical hypothesis classes will guide architecture design. Together, these theoretical advances will equip practitioners to choose ansätze and federation protocols that balance expressivity, convergence speed, and robust generalization on NISQ‐era devices. Further work could implement parallelization strategies, effcient qubit routing or dynamic circuit compilation (based on device’s states) to improve convergence.
@@ 8.3 Hardware Efficiency and Middleware Integration
Latency further increases challenges. As illustrated in Fig. 11, quantum middleware contains many points of bottlenecks and delays. Queue delays on public QPUs often span 10–60s per job, while compilation, including routing and gate‐reduction passes, can add 100–300ms per circuit [138, 130]. Heterogeneous client capabilities further distort global updates. Some nodes compile and execute circuits in under a second, others take tens of seconds, leading to stale gradients and model skew [131]. Meanwhile, real‐time error rates can drift by 5–10% over a single training day, mandating on‐the‐fly recalibration and continuous fidelity monitoring [76].
Looking forward, co‐design strategies seem promising to reduce these bottlenecks. Dynamic qubit remapping algorithms can reduce SWAP overhead by up to 25%, preserving circuit depth under tight connectivity constraints [139, 83]. Embedding error‐mitigation routines such as randomized compiling or zero‐noise extrapolation directly into the compilation pipeline, can recover 50–70% of lost fidelity without full quantum error correction [39]. From an architectural perspective, a federated hardware abstraction layer may be incorporated that could monitor live calibration metrics and queue statuses, enabling QFL frameworks to schedule gates and adjust ansatz depth automatically based on each client’s current noise profile [56]. These directions chart a course toward hardware‐aware QFL that adapts in real time, scales across diverse devices, and sustains federated convergence even as NISQ platforms evolve.
@@ 9 Insights to Experimental Evaluation
TABLE VI: Experimental details of recent and benchmark contributions.
Paper
@@
Accuracy Libraries
[49] Makecircles[140] - 100
@@
[77]
CIFAR-10[141],
Cats vs dogs[142]
100 100 FedAvg [1]
@@
Qulacs [143],
PyTorch [144],
PennyLane [145]
[50] Synthetic 30 - FedAvg [1] 98%
Cirq [146],
Tensorflow Quantum [147],
Tensorflow Federated [148]
[52] MNIST[149] 6 200 FedAvg [1] 99% -
[64] MNIST[149] - 20 - 81%
Tensorflow Federated [148],
Tensorflow Quantum [147]
[63]
@@
ECG dataset[150], and
Diabetes Health Indicator[151]
- - Consensus Average [152] - Tensor Toolbox (Matlab) [153]
[62]
RSNA chest X-ray[154],
NIH chest X-ray[155],
ADNI MRI-scan[156],
Kidney CT-scan[157]
- 35 FedAvg [1] 91%–98%
Tensorflow Quantum [147],
Tensor Network Library [158]
[61]
@@
5 15 FedAvg [1] - -
[60]
MNIST[149],
COVID-19[159]
@@
PyTorch [144],
PennyLane [145]
[19]
Iris[161] ,
Breast cancer, DNA[162]
@@
PennyLane [145],
PyTorch [144]
[89]
MedNIST[164],
COVID-19[159]
@@
FedAvg [1],
FedSGD[1]
@@
Tensorflow Quantum [147]
[103] Synthetic 5 - HQsFL[103] -
Scikit-Learn [165],
Tensorflow [166], PyTorch [144]
[104]
MNIST[149],
Moons[167], Synthetic
@@
PennyLane [145],
Scikit-Learn [165]
[105]
MNIST[149],
WDBC[168]
@@
PennyLane [145],
Qiskit [138]
[106] Cats vs Dogs[142] 100 - QFL-DP[169] 98%
PyTorch [144],
PyVacy [170]
[109]
MNIST[149],
CIFAR-10[141]
@@
[116] - - 150 FedAvg [1] -
@@
Cryptography [171],
PyCryptodome [172]
[117] MNIST[149] 5, 10, 15
@@
FedAvg [1] 85% -
[122] OASIS MRI[173] - - FedAvg [1] 78% -
[123] NSL-KDD[174] 5-25
@@
FedAvg [1] 98% PennyLane [145]
@@
Fatal Health[175],
Wisconsin Breast Cancer[176]
2, 4, 8 - FedAvg [1] 93%
PyTorch [144],
PennyLane [145]
@@
MNIST[149],
FMNIST[177]
- 100 FedAvg [1] 52% -
[131] mini-MNIST[178]
100 FedAvg [1] 52% -
@@
New England 39-bus[179] ,
Nordic system VSA[180]
6 - FedAvg [1] 98% -
[130]
CIFAR-10[141],
MNIST[149], FMNIST[177],
California Housing[181]
15 - FedAvg [1] 92%
@@
quantum annealer [182],
Gurobi LP solver [183]
[126]
CASIA[184],
RAVDESS[185], EMO-DB[186]
2, 4, 8 100 FedAvg [1] 74% PennyLane [145]
[125] NTUH[187] - -
@@
Weighted Averaging [188]
@@
Tensorflow Quantum [147],
Tensorflow Federated [148]
@@
GAIA Initiative[189]
@@
Weighted Averaging [188]
@@
OpenAI Gym [190],
Tensorflow Federated [148],
Tensorflow Quantum [147]

### v4

@@ 7.3 Quantum Secure Multi-Party Computation
In conventional FL systems, challenges such as gradient inversion attacks, which enable the reconstruction of private gradients, are prevalent. The authors in [QFL-GradientDP] addressed this issue by utilizing quantum states to encode gradients, thereby securing multi-party contributions to the global model from adversaries. The proposed framework incorporates protocols that use private inner-product estimation and incremental learning, collectively allowing contributions from each party without exposing individual gradients. In another work [QFL10], a quantum secure aggregation scheme is proposed to facilitate secure multi-party contributions by aggregating local models using quantum secret sharing, where participants encode model gradients into qubits and transmit them via Greenberger–Horne–Zeilinger (GHZ) states [greenberger2009ghz]. Quantum entanglement detects eavesdropping, ensuring data confidentiality. Further, in [QFL-HQK], a hybrid-quantum-key based secure FL system is designed for multi-party applications, particularly in the medical field. This system uses quantum key distribution to securely share parameters and hybrid quantum keys to protect data exchanges and model contributions from external attacks.
$\bullet$ Untrusted but non-adversarial computation nodes: Many QFL deployments involve service providers or benign participants that execute parts of the workflow but cannot be granted access to plaintext updates or model states, such as an outsourced PQC execution on a cloud QPU provider or aggregation nodes with stricter data-handling policies. These nodes need not necessarily be an attacker and thus must be carefully considered in the QFL framework.
A practical approach is to delegate quantum subroutines while concealing inputs and program structure from the compute node. It can be done by arranging the interaction so that an untrusted server performs the required quantum operations yet learns negligible information about the task beyond coarse resource usage. Such protocols have been proposed for low-latency, reliability-critical settings and are directly compatible with PQC calls within a QFL round [zaman2023concealed]. When QFL relies on intermediate routing or publication layers, anonymous transport primitives help prevent metadata leakage even if those layers are honest-but-curious. Recent results on controlled quantum anonymity  [khan2025controlledpub], quantum anonymous private information retrieval in distributed networks [khan2022qapir], and quantum anonymous collision detection for coordinating access [khan2021collision] provide building blocks for hiding the association between a client and its update/query while maintaining accountability and liveness in multi-party environments. In bandwidth-limited regimes, quantum semantic communication has also been explored to reduce exposure of raw symbols by transmitting meaning-level representations over shared links, which can lower leakage through intermediate nodes when combined with keying and aggregation [khalid2023qsemantic].
For QFL, these primitives map cleanly onto common roles. If the compute node is untrusted, clients keep data and classical parameters locally, delegate only the PQC evaluations with concealed inputs, and request verifiable transcripts of the delegated rounds for audit. If the aggregator is untrusted, clients retain their secure aggregation scheme and hybrid-keying and additionally route updates anonymously to hide which client contributed a given update. If aggregation nodes are untrusted, anonymous collision-detection and semantic communication reduce side-channel leakage while preserving throughput.
$\bullet$ Anonymous-network extensions: A distinct yet emerging strand of work complements quantum secure multi-party computation by focusing on anonymity within quantum networks. Mechanisms such as controlled quantum anonymous publication [khan2025controlledpub], quantum anonymous private information retrieval for distributed networks [khan2022qapir], and quantum anonymous collision detection for coordinating access to shared channels [khan2021collision] provide building blocks for concealing client identities and query patterns while maintaining verifiability. In parallel, concealed quantum telecomputation has been proposed for ultra-reliable low-latency communication contexts, enabling delegated computation on untrusted infrastructure without revealing inputs or program structure [zaman2023concealed]. Incorporating such primitives into QFL frameworks would extend beyond confidentiality of data values to protection of user metadata and communication traces, thereby strengthening privacy guarantees under both honest-but-curious and active adversarial models. This trajectory also opens avenues for future research on privacy-centric designs, where federated optimization is augmented with anonymity-preserving layers to reduce information leakage through routing and scheduling.
@@
Finally, Analyzing $\mathrm{QUE}$ for secure multi‐party computation schemes indicates near‐maximal qubit utilization across protocols. The inner‐product encoding in [QFL-GradientDP] uses a four‐qubit, three‐layer circuit ($\mathrm{Q}_{alloc}=3$, $L=3$) obtaining $\mathrm{QUE}=1.0$ and thus adopts synchronous aggregation to leverage full‐batch quantum metric estimation; the GHZ‐based secret sharing in [QFL10] runs on a three‐qubit, two‐layer design ($\mathrm{Q}_{alloc}=3$, $L=2$) with $\mathrm{QUE}=1.0$ and similarly favors synchronous updates to maintain entanglement integrity; and the quantum‐key DFL in [QFL-HQK] employs a five‐qubit, four‐layer ansatz ($\mathrm{Q}_{alloc}=5$, $L=4$) to achieve a high $\mathrm{QUE}\approx 1.0$ paired with hybrid classical–quantum fusion to balance key‐distribution overhead with model convergence. Despite their efficient parameter‐gate ratios and aligned model strategy choices, these multi-party computation frameworks have significant limitations. Firstly, they assume ideal entanglement distribution and noise‐free communication, omit analysis under non‐IID data splits, and lack adaptive mechanisms for client dropout or key‐renewal scheduling. Future work should explore decoherence‐aware aggregation protocols, resilient key management strategies in dynamic networks, and benchmark suites that jointly measure privacy, QUE, and end‐to‐end latency in dynamic QFL environments.
@@ 8 Miscellaneous
This section aims to elaborate on the practical usage and implementation of QFL, while discussing studies that apply it to real-life datasets that demonstrate QFL’s true potential. It also discusses the trainability and learnability aspects of QFL to further discuss challenges and opportunities.
@@ 8.1 Real-life Applications of QFL
Numerous studies have also explored QFL to enhance performance and address the computational demands in diverse application domains. Fig. 11 shows application domains where QFL research has shown its footprints.
In healthcare contexts, federated quantum designs often emphasize data confidentiality. One dementia-classification method [tanbhir2025quantuminspiredprivacypreservingfederatedlearning] employs QKD for encrypting parameter updates and safeguarding sensitive MRI data across hospital nodes. Despite a 1% increment in computational overhead, classification accuracy matched the baseline CNN (roughly 94%) while guaranteeing that eavesdroppers cannot easily invert gradients. Another proposal targets secure intrusion detection in consumer networks [10679217], achieving up to 3–4% gains in detection rates on complex traffic features when harnessing small-scale QNNs and homomorphic gradient aggregations. Addressing data heterogeneity is a central challenge in large-scale QFL. The Dynamic Aggregation QFL (DAQFL) approach [DAQFL] introduces an adaptive weighting scheme based on local accuracy improvements. In healthcare IoT benchmarks, DAQFL achieves a 2–6% accuracy gain over baseline quantum federated averaging when dealing with skewed or long-tailed data distributions. QFL has also shown efficacy for sensitive clinical tasks such as pain-level detection. In [balasubramani2025novel], ECG signals are processed with a quantum-transfer-learning pipeline, integrated within a federated scheme to preserve patient privacy. Compared to conventional CNN-based methods, the quantum model improved accuracy by nearly 2%, reaching approximately 94.8% for multi-tier pain labeling.
@@
Moving into automotive applications, QFSM [qu2024qfsm] addresses speech-emotion recognition in vehicular networks by employing quantum minimal-gated recurrent units. This design shortens the gating operations and cuts training convergence time by roughly 10–15%. Field tests in a 5G IoV context report up to 2% higher accuracy than classical gated recurrent units on large-scale SER corpora, while also offering improved noise robustness in fluctuating channel conditions. Vehicular networks have also been extended into metaverse-based coordination. A quantum-enhanced FL framework [10758814] uses hierarchical clustering of vehicular nodes, where QNN states are compressed via quantum principal-component analysis. Adaptive reinforcement scheduling then modulates local updates, ensuring resource balancing across immersive metaverse tasks. Simulation results reflect around 15% lower communication overhead compared to standard FL. A related thread investigates quantum FL in edge or drone-assisted scenarios. One surveillance solution [10829961] allows UAVs to transmit partial circuit parameters, dynamically adjusting QNN layer depths based on instantaneous channel quality. This reduces the communication load by nearly 20% while retaining top-1 accuracy comparable to a static-depth model on an aerial license-plate dataset.
Another notable approach, ESQFL [10734228], integrates digital twin frameworks for real-time power-grid stability analysis. By linking quantum FL updates to digital replicas of physical bus nodes, the global model can detect incipient voltage collapse and dynamically schedule corrective actions. Shapley value calculations explain each node’s contribution in an interpretable manner. Grid simulations showed a 4% improvement in detecting instability events and a 10% boost in overall reliability.
Beyond the model-training process, some works explore hybrid optimization for FL resource management. One Benders’ decomposition algorithm [10636852] splits scheduling of workers and continuous learning-rate decisions into subproblems, solved alternately on a quantum annealer and a classical solver. At moderate scales, the method reduces runtime by up to 70% versus purely classical MILP solutions, indicating promise for large-scale FL scheduling. A further refinement involves quantum gate customization. Entanglement-controlled QFL [10878965] employs specialized universal gates to fine-tune the entanglement entropy in multi-depth QNNs. This strategy mitigates the barren-plateau problem by balancing inter-qubit interference. Experiments indicate consistent convergence on non-iid data with minimal gradient vanishing.
@@ 8.2 Trainability and Learnability of QFL
Federated quantum models combine the expressive power of parameterized quantum circuits with the decentralized optimization of FL. Yet they inherit two intertwined theoretical challenges: ensuring effective optimization (trainability) and guaranteeing robust generalization (learnability).
As circuit depth $L$ or qubit count $Q$ increases, gradient magnitudes can vanish exponentially, a phenomenon known as the barren plateau [mcclean2018barren, Cerezo2021]. In QFL, each client’s local update may fall below machine noise, forcing the server to perform many more communication rounds to recover signal [QFL-QNGD]. Empirical results on hardware-efficient ansatz show that limiting depth to $L\leq 4$ sustains gradients above numeric precision for up to 16 qubits, reducing required rounds by 30–50% compared to deeper designs [abbas2021power]. Quantifying convergence, recent studies measure the number of gradient‐descent steps or federated rounds needed to reach a target accuracy. For example, FQNGD achieves 90% of its asymptotic accuracy within 20 rounds on standard benchmarks, compared to 40 rounds for FedAvg with the same ansatz [QFL-QNGD]. This halving of communication cost stems from quantum natural gradient descent, which preconditions updates by the inverse quantum Fisher information matrix, enhancing per‐round progress up to 2× on ideal simulators [QFL-QNGD]. Additional acceleration techniques further improve trainability. Layerwise training optimizes one ansatz block at a time and preserves the gradient signal by avoiding global barren plateaus. Alternatively, smart initialization reuses parameters from smaller or related circuits to jump‐start optimization, cutting VQA iterations by up to 60% in QAOA settings [ML-Optimization, skolik2023quantum]. Together, these methods yield hybrid protocols that converge in fewer than 25 rounds even under moderate noise, making QFL practical on NISQ hardware.
On the other hand, learnability examines how well a QFL model trained on local data samples extends to unseen inputs. Classical FL theory derives generalization bounds via VC‐dimension and algorithmic stability; quantum analogues are emerging. Shallow, low‐entanglement circuits behave like kernel machines with controlled Rademacher complexity, suggesting that $\mathcal{O}(\log N)$ qubits suffice to generalize on $N$ training examples [schuld2021effect]. Yet over‐parameterized quantum models exhibit a double descent curve i.e., after an interpolation threshold, additional parameters can improve generalization under appropriate regularization [Caro_2022, PRXQuantum.2.040321].
@@
where $n_{\rm classical}$, $n_{\rm quantum}$ are local sample sizes and $C_{\rm classical},C_{\rm quantum}$ depend on circuit expressivity [bu2021rademacher]. Stability‐based bounds further relate the sensitivity of the global model to perturbations in client updates, with quantum algorithmic stability results implying that small‐norm updates preserve generalization under noisy mixing [yang2025stabilitygeneralizationquantumneural].
Fully characterizing the federated barren‐plateau frontier requires mapping how depth $L$, width $Q$, and client heterogeneity jointly determine gradient norms and convergence rates. On the learnability side, deriving optimal VC‐dimension and Rademacher‐complexity bounds for hybrid quantum–classical hypothesis classes would guide architecture design. Further work could implement parallelization strategies, efficient qubit routing, or dynamic circuit compilation (based on the device’s states) to improve convergence.
@@ 8.3 Hardware Efficiency and Middleware Integration
As illustrated in Fig. 12, quantum middleware contains many points of bottlenecks and delays. Queue delays on public QPUs often span 10–60s per job, while compilation, including routing and gate‐reduction passes, can add 100–300ms per circuit [qiskit, 10636852]. Heterogeneous client capabilities further distort global updates. Some nodes compile and execute circuits in under a second, others take tens of seconds, leading to stale gradients and model skew [10878965]. Meanwhile, real‐time error rates can drift by 5–10% over a single training day, mandating on‐the‐fly recalibration and fidelity monitoring [error1].
Looking forward, co‐design strategies seem promising to reduce these bottlenecks. Dynamic qubit remapping algorithms can reduce SWAP overhead by up to 25%, preserving circuit depth under tight connectivity constraints [khandavilli2023towards, nash2020quantum]. Embedding error‐mitigation routines such as randomized compiling or zero‐noise extrapolation directly into the compilation pipeline, can recover 50–70% of lost fidelity without full quantum error correction [Cerezo2021]. From an architectural perspective, a federated hardware abstraction layer may be incorporated that could monitor live calibration metrics and queue statuses, enabling QFL frameworks to schedule gates and adjust ansatz depth automatically based on each client’s current noise profile [ferrari2021compiler]. These directions chart a course toward hardware‐aware QFL that adapts in real time, scales across diverse devices, and sustains federated convergence even as NISQ platforms evolve.
@@ 9 Insights to Experimental Evaluation
TABLE VII: Experimental details of recent and benchmark contributions.
[…]
@@
[…]
[QFL-TL] Makecircles[makecircles_sklearn] - 100
@@
[QFL-TL-2]
CIFAR-10[krizhevsky2009learning],
Cats vs dogs[cats_vs_dogs_kaggle]
100 100 FedAvg [mcmahan2017communication]
@@
Qulacs [qulacs],
PyTorch [pytorch],
PennyLane [pennylane]
[FQE1] Synthetic 30 - FedAvg [mcmahan2017communication] 98%
Cirq [cirq],
Tensorflow Quantum [tfq],
Tensorflow Federated [tff]
[QFL-QNGD] MNIST[lecun1998gradient] 6 200 FedAvg [mcmahan2017communication] 99% -
[QFL-QAOA] MNIST[lecun1998gradient] - 20 - 81%
Tensorflow Federated [tff],
Tensorflow Quantum [tfq]
[QFL-Tensor2]
@@
ECG dataset[mitbih_arrhythmia], and
Diabetes Health Indicator[pima_indians_diabetes_uci]
- - Consensus Average [olfati2007consensus] - Tensor Toolbox (Matlab) [tensorToolbox]
[QFL-Tensor]
RSNA chest X-ray[rsna_chest_xray_pneumonia],
NIH chest X-ray[wang2017chestxray8],
ADNI MRI-scan[adni],
Kidney CT-scan[kits19_data]
- 35 FedAvg [mcmahan2017communication] 91%–98%
Tensorflow Quantum [tfq],
Tensor Network Library [tensornetwork]
[QFL-QLSTM]
@@
5 15 FedAvg [mcmahan2017communication] - -
[QFL-QFuzzyNN]
MNIST[lecun1998gradient],
COVID-19[covid_chestxray_dataset]
@@
PyTorch [pytorch],
PennyLane [pennylane]
[QFL2]
Iris[fisher1936iris] ,
Breast cancer, DNA[dna_splice_junction_uci]
@@
PennyLane [pennylane],
PyTorch [pytorch]
[QFL-QCNN]
MedNIST[medmnistv2],
COVID-19[covid_chestxray_dataset]
@@
FedAvg [mcmahan2017communication],
FedSGD[mcmahan2017communication]
@@
Tensorflow Quantum [tfq]
[QFL-DPFHE] Synthetic 5 - HQsFL[QFL-DPFHE] -
Scikit-Learn [scikit-learn],
Tensorflow [tensorflow], PyTorch [pytorch]
[QFL-DPNoise]
MNIST[lecun1998gradient],
Moons[make_moons_sklearn], Synthetic
@@
PennyLane [pennylane],
Scikit-Learn [scikit-learn]
[QFL-BlindC]
MNIST[lecun1998gradient],
WDBC[wdbc_uci]
@@
PennyLane [pennylane],
Qiskit [qiskit]
[QFL-DP] Cats vs Dogs[cats_vs_dogs_kaggle] 100 - QFL-DP[qfldp2021] 98%
PyTorch [pytorch],
PyVacy [pyvacy]
[QFL10]
MNIST[lecun1998gradient],
CIFAR-10[krizhevsky2009learning]
@@
[gharavi2025pqbflpostquantumblockchainbasedprotocol] - - 150 FedAvg [mcmahan2017communication] -
@@
Cryptography [postquantum],
PyCryptodome [pycryptodome]
[maouaki2025qfalquantumfederatedadversarial] MNIST[lecun1998gradient] 5, 10, 15
@@
FedAvg [mcmahan2017communication] 85% -
[tanbhir2025quantuminspiredprivacypreservingfederatedlearning] OASIS MRI[oasis_mri_marcus2007] - - FedAvg [mcmahan2017communication] 78% -
[10679217] NSL-KDD[NSL-KDD] 5-25
@@
FedAvg [mcmahan2017communication] 98% PennyLane [pennylane]
[DAQFL]
Fatal Health[fetal_health],
Wisconsin Breast Cancer[breast_cancer_wisconsin_diagnostic]
2, 4, 8 - FedAvg [mcmahan2017communication] 93%
PyTorch [pytorch],
PennyLane [pennylane]
@@
MNIST[lecun1998gradient],
FMNIST[fashion_mnist_xiao2017]
- 100 FedAvg [mcmahan2017communication] 52% -
[10878965] mini-MNIST[mini_mnist_no_official]
100 FedAvg [mcmahan2017communication] 52% -
@@
New England 39-bus[new_england_39bus_athay1979] ,
Nordic system VSA[nordic_test_system_cigre]
6 - FedAvg [mcmahan2017communication] 98% -
@@
CIFAR-10[krizhevsky2009learning],
MNIST[lecun1998gradient], FMNIST[fashion_mnist_xiao2017],
California Housing[cal_housing_sklearn]
15 - FedAvg [mcmahan2017communication] 92%
@@
quantum annealer [dwave],
Gurobi LP solver [gurobi]
[qu2024qfsm]
CASIA[casia_iris],
RAVDESS[livingstone2018ravdess], EMO-DB[emo_db_berlin]
2, 4, 8 100 FedAvg [mcmahan2017communication] 74% PennyLane [pennylane]
[balasubramani2025novel] NTUH[ntuh_ecg_physionet] - -
@@
Weighted Averaging [Bonawitz2017SecAgg]
@@
Tensorflow Quantum [tfq],
Tensorflow Federated [tff]
@@
GAIA Initiative[didi_gaia]
@@
Weighted Averaging [Bonawitz2017SecAgg]
@@
OpenAI Gym [openai_gym],
Tensorflow Federated [tff],
Tensorflow Quantum [tfq]